OpenStudy (torii897):

pleaaseeeee help The expression 2(l + w) is used to calculate the perimeter of a rectangle, where l is length and w is width. If the length is Fraction 2 over 3 unit and the width is Fraction 1 over 3 unit, what is the perimeter of the rectangle in units? Fraction A:2 over 3 unit B:1 unit c:1Fraction 2 over 3 units D: 2 units

OpenStudy (torii897):

@LynFran

OpenStudy (torii897):

@Michele_Laino ????

OpenStudy (michele_laino):

we have: \[\Large perimeter = 2\left( {l + w} \right) = 2\left( {\frac{2}{3} + \frac{1}{3}} \right) = ...?\]

OpenStudy (michele_laino):

hint: \[\Large \frac{2}{3} + \frac{1}{3} = \frac{{2 + 1}}{3} = ...?\]